The television show CSI: Shoboygan has been successful for many years. That show recently had a share of 18, meaning that among the TV sets in use, 18% were tuned to CSI: Shoboygan. Assume that an advertiser wants to verify that 18% share value by conducting its own survey, and a pilot survey begins with 14 households have TV sets in use at the time of a CSI: Shoboygan broadcast. Find the probability that none of the households are tuned to CSI: Shoboygan. P(none) = Find the probability that at least one household is tuned to CSI: Shoboygan. P(at least one) = Find the probability that at most one household is tuned to CSI: Shoboygan. P(at most one) = If at most one household is tuned to CSI: Shoboygan, does it appear that the 18% share value is wrong? (Hint: Is the occurrence of at most one household tuned to CSI: Shoboygan unusual?)
Answer #1

X ~ Binomial (n,p)

Where n = 14, p = 0.18

Binomial probability distribution is

P(X) = nCx px ( 1 - p)n-x

a)

P( X = 0) = ( 1 - 0.18)14

= 0.0621

b)

P( X >= 1) = 1 - P( X = 0)

= 1 - 0.0621

= 0.9379

c)

P( X <= 1) = P(X = 0) + P( X = 1)

= ( 1 - 0.18)14 + 0.18 * ( 1 - 0.18)13

= 0.2531

Since 0.2531 is greater than 0.05, the event is not unusual.